OpenStudy (anonymous):

One of the intersecting line angles is 8 times bigger than the other angle. Find the smaller angle.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

A1= Angle 1 A2= Angle 2 A1+A2=180 A2=8*A1 A1+8*A1=180 9*A1=180 A1=20

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

I do. Thanks!